# Bonsai: Gradient-free Graph Distillation for Node Classification

### v1

@@ Abstract
Graph distillation has emerged as a promising avenue to enable scalable training of Gnns by compressing the training dataset while preserving essential graph characteristics. Our study uncovers significant shortcomings in current graph distillation techniques. First, the majority of the algorithms paradoxically require training on the full dataset to perform distillation. Second, due to their gradient-emulating approach, these methods require fresh distillation for any change in hyperparameters or Gnn architecture, limiting their flexibility and reusability. Finally, they fail to achieve substantial size reduction due to synthesizing fully-connected, edge-weighted graphs. To address these challenges, we present Bonsai, a novel graph distillation method empowered by the observation that computation trees form the fundamental processing units of message-passing Gnns. Bonsai distills datasets by encoding a careful selection of exemplar trees that maximize the representation of all computation trees in the training set. This unique approach imparts Bonsai as the first linear-time, model-agnostic graph distillation algorithm for node classification that outperforms existing baselines across $6$ real-world datasets on accuracy, while being $22$ times faster on average. Bonsai is grounded in rigorous mathematical guarantees on the adopted approximation strategies making it robust to Gnn architectures, datasets, and parameters.
@@ 1 Introduction and Related Works
Graph Neural Networks (Gnns) have shown remarkable success in various predictive tasks on graph data (Veličković et al., 2018; Kipf & Welling, 2017; Hamilton et al., 2017). However, real world graphs often contain millions of nodes and edges making the training pipeline slow and computationally demanding (Hu et al., 2021). This limitation hinders their adoption in resource-constrained environments (Miao et al., 2021) and applications dealing with massive datasets (Hu et al., 2021). Graph distillation (also called condensation) has emerged as a promising way to bypass this bottleneck (Jin et al., 2021; Liu et al., 2024a; b; Zheng et al., 2023). The objective in graph distillation is to synthesize a significantly smaller distilled data set that retains the essential information of the original data. By training Gnns on these distilled datasets, we can achieve comparable performance while reducing computational overhead and storage costs. This makes Gnns more accessible and practical for a wider range of applications, including those with limited computational resources and large-scale datasets. In this work, we study the problem of graph distillation for node classification.
@@ 1.1 Existing works and their Limitations
Table 1 presents existing graph distillation algorithms proposed for node classification and their characterization across various dimensions. We omit listing DosCond (Jin et al., 2022), Mirage (Gupta et al., 2024), and KiDD (Xu et al., 2023) in Table 1 since they are designed for graph classification, whereas we focus on node classification. The visual representation in Table 1 allows for quick identification of strengths and limitations of existing methods. We elaborate these limitations.
$\bullet$ Full Gnn training is a pre-requisite: The fundamental requirement of data distillation is that it should require less computational resources and time than training on the full dataset. However, this basic premise is violated by majority of the techniques (See Table 1) since they adopt a design where training the target Gnn on the full trainset is a prerequisite to distillation. These algorithms adopt a gradient-dependent optimization framework. They first train a Gnn on the full trainset and extract gradients of model parameters over epochs (Sfgc and Gcsr require training on the full dataset up to 200 and 100 times, respectively). The distillation process is formulated as an optimization problem to create a distilled dataset that replicates the gradient trajectory observed in the original training set. This need for full dataset training for distillation contradicts its fundamental premise, which we address in our work.
Table 1: Characterization of existing graph distillation algorithms. Cells shaded in Red indicate the presence of an undesirable property, while Green represents their absence.
Algorithm Requires training Gnn on full-dataset Generates fully-connected graph Model-specific distillation
@@
Table 2: We present the number of edges in the distilled graphs produced by GCond (Jin et al., 2021) and compare them to the full dataset. Red cells indicate cases where the distilled graph have more edges than the full dataset. The indicated node ratios are taken from the values used in GCond. Since all other baselines also produce fully-connected graphs, the numbers from this table translate to them as well.
Dataset Distilled Full Dataset
($r_{n}$ or node ratio) edge ratio #edges #edges
Cora (5.2%) 186% 19,600 10,556
Citeseer (3.6%) 158% 14,400 9,104
Pubmed (0.3%) 4% 3,600 88,648
Flickr (1%) 88% 795,664 899,756
Ogbn-arxiv (0.5%) 31% 715,716 2,315,598
Reddit (0.2%) 1% 216,225 23,213,838
$\bullet$ Distilling to a fully-connected graph: In a message-passing Gnn, the computation cost of each forward pass is $\mathcal{O}(|\mathcal{E}|)$, where $\mathcal{E}$ denotes the set of edges. Consequently, the computational effectiveness of graph distillation is primarily determined by the reduction in edge count between the original and distilled graphs, rather than node count alone. However, current graph distillation algorithms (see Table 1) quantify the condensation ratio based on the node count. Specifically, given a compression ratio $r_{n}$, it synthesizes a weighted, fully-connected dense adjacency matrix for the distilled graph of the size $\frac{|\mathcal{V}|}{r_{n}}\times\frac{|\mathcal{V}|}{r_{n}}$, where $\mathcal{V}$ denotes the node set. This design, which is inconsistent with the computation structure of Gnns, can lead to a distilled graph with minimal reduction in edge count and, in some cases, may even result in a graph with more edges than the original dataset (See Table 2).
$\bullet$ Model-specific distillation: Gradients of model weights are influenced by the specific Gnn architecture and hyper-parameters (e.g., number of layers, hidden dimensions, dropout rates). Consequently, any architectural change, such as switching from a Gcn (Kipf & Welling, 2017) to a Gat (Veličković et al., 2018), or hyper-parameter adjustments, necessitates a new round of distillation.
@@ 1.2 Contributions
To address the limitations outlined in Table 1, we present Bonsai  [^1] Inspired by the art of Bonsai, which transforms large trees into miniature forms while preserving their essence, our graph distillation algorithm gracefully prunes redundant computation trees, creating a distilled graph that is significantly smaller yet maintains comparable performance..
  • Gradient-free distillation: Instead of replicating the gradient trajectory, Bonsai emulates the distribution of input data processed by message-passing Gnns. By shifting the computation task to the pre-learning phase, Bonsai achieves independence from hyper-parameters and model architectures as long as it adheres to a message-passing Gnn framework like Gat (Veličković et al., 2018), Gcn (Kipf & Welling, 2017), GraphSage (Hamilton et al., 2017), Gin (Xu et al., 2019), etc. Moreover, this addresses a critical limitation of existing graph distillation algorithms that necessitates training on the entire training dataset.
  • Novel algorithm design: Bonsai is empowered by the observation that any message-passing Gnn decomposes a graph of $n$ nodes into $n$ rooted computation trees. Furthermore, topologically similar computation trees generate similar embeddings regardless of the Gnn being used (Xu et al., 2019; Togninalli et al., 2019). Bonsai exploits this observation to identify a small subset of diverse computation trees, called exemplars, that are located in dense regions and thereby representative of the full set. Hence, the induced subgraph spanned by the exemplars forms an effective distilled set.
  • Empirical evaluation: We perform rigorous benchmarking incorporating state-of-the-art graph distillation algorithms on $6$ real-world datasets containing up to hundreds of millions of edges. Our analysis establishes that Bonsai (1) achieves higher prediction accuracy, (2) produces $22$ times faster distillation times despite being CPU-bound in contrast to GPU-bound distillation of baselines, and (3) exhibits superior robustness to Gnn architectures and datasets.
@@ 2 Problem Formulation and Preliminaries
The problem of graph distillation for node classification is defined as follows.
Problem 1 (Graph Distillation).
@@ 2.1 Computation structure of Gnns
A computation tree, also known as a receptive field, describes how information propagates through the graph during the neural network’s operation (Xu et al., 2019). Formally, it is defined as follows.
@@
If $\mathcal{T}_{v}^{L}$ is isomorphic to $\mathcal{T}_{u}^{L}$, then $\mathbf{h}_{v}^{L}=\mathbf{h}_{u}^{L}$. This follows from the property that the expressive power of a message-passing Gnn is upper bounded by the Weisfeiler-Lehman test (1-WL) (Xu et al., 2019). This implies that if the $L$-hop neighborhoods of two nodes are indistinguishable by 1-WL, then their representations will be the same. The 1-WL test cannot differentiate between isomorphic computation trees (Shervashidze et al., 2011).
@@
To test our hypothesis, in Fig. 2, we examine all node pairs within a specified distance threshold based on the Weisfeiler-Lehman (WL) kernel. We then analyze the correlation between their unsupervised WL-distance and the $L_{2}$ distance between their Gnn embeddings (Gcn with $2$ layers, trained with cross-entropy loss). The results reveal a compelling trend: as we decrease the distance threshold, we observe a strengthening correlation between WL-distance and Gnn embedding distance. This pattern strongly supports our hypothesis, indicating that computation trees that are proximate in the WL-space also exhibit proximity in the Gnn embedding space.
@@ 3 Bonsai: Proposed Methodology
    Similar Gnn embeddings generate comparable outputs, resulting in similar impacts on the loss function and, consequently, on the gradients.
Figure 3: Pipeline of the proposed algorithm for graph distillation
@@
  • Representativeness: Each exemplar should be similar to a large number of other computation trees from the input training set, ensuring it captures common structural patterns. We quantify the representative power of a computation tree using the idea of reverse $k$ nearest neighbors (§ 3.1).
  • Diversity: The set of exemplars should be diverse, maximizing the coverage of different structural patterns present in the original graph (§ 3.2).
By optimizing for both representativeness and diversity, Bonsai ensures that the union of computation trees similar to the chosen exemplars encompasses a broad and representative spectrum of the original graph’s structural properties. The induced subgraph spanned by the nodes in these exemplar trees forms the initial distilled set. This initial version undergoes further refinement through an iterative process of edge sparsification and enrichment to reduce $O(|\mathcal{E}|)$ complexity of Gnns.
@@ 3.1.1 Sampling for Scalable Computation of Reverse $k$-NN
Proof. For the formal proof, see App. A.1. Intuitively, if a tree participates in the $k$-NN of a large percentage of trees from the full dataset, then it will continue to do so in the sampled set as well. Lemma 1 establishes this formally resulting in the following positive implications.
@@ 3.2 Coverage Maximization
The exemplars, $\mathbb{A}_{greedy}$, selected by Alg. 1 provides an $1-1/e$ approximation, i.e., $\Pi(\mathbb{A}_{greedy})\geq\left(1-\frac{1}{e}\right)\Pi(\mathbb{A}^{*})$,
@@
Initial distilled graph: The initial distilled graph $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\bm{X}_{s})$ is formed by extracting the induced subgraph spanned by the exemplar computation trees, i.e., $\mathcal{V}_{s}=\left\{u\mid\exists\mathcal{T}_{v}^{L}\in\mathbb{A}_{greedy},\>u\in\mathcal{N}_{L}(v)\right\}$, $\mathcal{E}_{s}=\{(u,v)\mid u\in\mathcal{V}_{s},v\in\mathcal{V}_{s}\}$ and $\bm{X}_{s}=\{\mathbf{x}_{v})\mid v\in\mathcal{V}_{s}\}$. Set $\mathcal{N}_{L}(v)$ contains nodes within $L$ hops from $v$.
3.3 Sparsification and Enrichment
Recall that the complexity of a Gnn forward pass is linear to the number of edges in the graph. In the final step, we seek to sparsify the graph induced by the exemplar trees. Furthermore, in the additional space created due to sparsfication, we include more exemplar trees, resulting in further magnification of the representative power. This process is formalized as follows.
Let $\mathcal{V}_{S}$ be the node set in the initial distilled graph, which can be partitioned into two disjoint subsets: $\mathcal{V}_{root}$ and $\mathcal{V}_{ego}$, where $\mathcal{V}_{root}=\{v\mid\exists\mathcal{T}_{v}^{L}\in\mathbb{A}\}$, the root nodes, that serve as the root node of some exemplar in $\mathbb{A}_{greedy}$, and $\mathcal{V}_{ego}=\mathcal{V}_{S}\setminus\mathcal{V}_{root}$, the ego nodes, that are included because they fall within the $L$-hop neighborhood of a root node. We propose to employ Personalized PageRank (PPR) (Page et al., 1998) with teleportation only to root nodes to identify and prune ego nodes with minimal impact on the root node embeddings. Starting with timestamp $t=0$, we proceed as follows:
  1. 1.
    Compute the PPR distribution $\bm{\pi}_{t}$ for $t\geq 0$ defined by:
    $\bm{\pi}_{t}=(1-\beta)\mathbf{A}\bm{\pi}_{t-1}+\beta\mathbf{e}$ (8)
  2. 2.
  3. 3.
  4. 4.
    Define the knee point $i_{knee}$ as:
  5. 5.
    Prune all ego nodes with PPR scores below $\bm{\pi}[i_{knee}]$.
  6. 6.
  7. 7.

### v3

# Bonsai: Gradient-free Graph Condensation for Node Classification

@@ Abstract
Graph condensation has emerged as a promising avenue to enable scalable training of Gnns by compressing the training dataset while preserving essential graph characteristics. Our study uncovers significant shortcomings in current graph condensation techniques. First, the majority of the algorithms paradoxically require training on the full dataset to perform condensation. Second, due to their gradient-emulating approach, these methods require fresh condensation for any change in hyper-parameters or Gnn architecture, limiting their flexibility and reusability. To address these challenges, we present Bonsai, a novel graph condensation method empowered by the observation that computation trees form the fundamental processing units of message-passing Gnns. Bonsai condenses datasets by encoding a careful selection of exemplar trees that maximize the representation of all computation trees in the training set. This unique approach imparts Bonsai as the first linear-time, model-agnostic graph condensation algorithm for node classification that outperforms existing baselines across $7$ real-world datasets on accuracy, while being $22$ times faster on average. Bonsai is grounded in rigorous mathematical guarantees on the adopted approximation strategies, making it robust to Gnn architectures, datasets, and parameters.
@@ 1 Introduction and Related Works
Graph Neural Networks (Gnns) have shown remarkable success in various predictive tasks on graph data such as node classification and link prediction (Veličković et al., 2018; Kipf & Welling, 2017; Hamilton et al., 2017; Nishad et al., 2021), modeling physical systems (Bhattoo et al., 2022; Thangamuthu et al., 2022; Bishnoi et al., 2023; 2024), and spatio-temporal data (Jain et al., 2021; Gupta et al., 2023). However, real-world graphs often contain millions of nodes and edges making the training pipeline slow and computationally demanding (Hu et al., 2021). This limitation hinders their adoption in resource-constrained environments (Miao et al., 2021) and applications dealing with massive datasets (Hu et al., 2021). Graph condensation (also called graph distillation) has emerged as a promising way to bypass this bottleneck (Jin et al., 2021; Liu et al., 2024a; b; Zheng et al., 2023). The objective in graph condensation is to synthesize a significantly smaller condensed data set that retains the essential information of the original data. By training Gnns on these condensed datasets, we can achieve comparable performance while reducing computational overhead and storage costs. This makes Gnns more accessible and practical for a wider range of applications, including those with limited computational resources and large-scale datasets. In this work, we study the problem of graph condensation for node classification.
@@ 1.1 Existing works and their Limitations
Table 1 presents existing graph condensation algorithms proposed for node classification and their characterization across various dimensions. We omit listing DosCond (Jin et al., 2022), Mirage (Gupta et al., 2024), and KiDD (Xu et al., 2023) in Table 1 since they are designed for graph classification, whereas we focus on node classification.
$\bullet$ Full Gnn training is a pre-requisite: The fundamental requirement of data condensation is that it should require less computational resources and time than training on the full dataset. However, this basic premise is violated by majority of the techniques (See Table 1) since they adopt a design where training the target Gnn on the full trainset is a prerequisite to condensation. These algorithms adopt a gradient-dependent optimization framework. They first train a Gnn on the full trainset and extract gradients of model parameters over epochs. The condensation process is formulated as an optimization problem to create a condensed dataset that replicates the gradient trajectory observed in the original training set. This need for full dataset training for condensation contradicts its fundamental premise, which we address in our work.
Table 1: Characterization of existing graph condensation algorithms. Cells shaded in Red indicate the presence of an undesirable property, while Green represents their absence.
Algorithm Requires training Gnn on full-dataset Condenses to a fully-connected graph [^1] Some algorithms sparsify the fully-connected graph based on edge weights. But this sparsification process requires training on the fully connected graph itself to identify the pruning threshold. Model-specific condensation
@@
GC-Sntk (Wang et al., 2024)
Exgc (Fang et al., 2024)
Geom (Zhang et al., 2024)
@@
$\bullet$ Node compression vs. Edge complexity: In a message-passing Gnn, the computation cost of each forward pass is $\mathcal{O}(|\mathcal{E}|)$, where $\mathcal{E}$ denotes the set of edges. Consequently, the computational effectiveness of graph condensation is primarily determined by the reduction in edge count between the original and condensed graphs, rather than node count alone. However, current graph condensation algorithms (see Table 1) quantify the condensation ratio based on the node count. This creates a fundamental disconnect between the metric used to evaluate compression (node count) and the computational goal of reducing $\mathcal{O}(|\mathcal{E}|)$, which directly governs training and inference time. Consequently, this misalignment can lead to a condensed graph that is less efficient than the original.
$\bullet$ Model-specific condensation: Gradients of model weights are influenced by the specific Gnn architecture and hyper-parameters (e.g., number of layers, hidden dimensions, etc.). Consequently, any architectural change, such as switching from a Gcn (Kipf & Welling, 2017) to a Gat (Veličković et al., 2018), or hyper-parameter adjustments, necessitates a new round of condensation.
@@ 1.2 Contributions
To address the limitations outlined in Table 1, we present Bonsai  [^2] Inspired by the art of Bonsai, which transforms large trees into miniature forms while preserving their essence, our graph condensation algorithm gracefully prunes redundant computation trees, creating a condensed graph that is significantly smaller yet maintains comparable performance..
  • Gradient-free condensation: Instead of replicating the gradient trajectory, Bonsai emulates the distribution of input data processed by message-passing Gnns. By shifting the computation task to the pre-learning phase, Bonsai achieves independence from hyper-parameters and model architectures as long as it adheres to a message-passing Gnn framework like Gat, Gcn, GraphSage, Gin, etc. Moreover, this addresses a critical limitation of existing graph condensation algorithms that necessitates training on the entire training dataset.
  • Novel algorithm design: Bonsai is empowered by the observation that any message-passing Gnn decomposes a graph of $n$ nodes into $n$ rooted computation trees. Furthermore, topologically similar computation trees generate similar embeddings regardless of the Gnn being used (Xu et al., 2019; Togninalli et al., 2019). Bonsai exploits this observation to identify a small subset of diverse computation trees, called exemplars, that are located in dense regions and thereby representative of the full set. Hence, the induced subgraph spanned by the exemplars forms an effective condensed set.
  • Empirical evaluation: We perform rigorous benchmarking incorporating state-of-the-art graph condensation algorithms on $7$ real-world datasets containing up to hundreds of millions of edges. Our analysis establishes that Bonsai (1) achieves higher prediction accuracy, (2) produces at least $7$-times faster condensation times despite being CPU-bound in contrast to GPU-bound condensation of baselines, and (3) exhibits superior robustness to Gnn architectures and datasets.
@@ 2 Problem Formulation and Preliminaries
The problem of graph condensation for node classification is defined as follows.
Problem 1 (Graph Condensation).
@@ 2.1 Computation structure of Gnns
A computation tree describes how information propagates through the graph during the neural network’s operation (Xu et al., 2019).
@@
If $\mathcal{T}_{v}^{L}$ is isomorphic to $\mathcal{T}_{u}^{L}$, then $\mathbf{h}_{v}^{L}=\mathbf{h}_{u}^{L}$ since the expressive power of a message-passing Gnn is upper bounded by the Weisfeiler-Lehman test (1-WL) (Xu et al., 2019).
@@
To test our hypothesis, in Fig. 2, we examine all node pairs within a specified distance threshold based on the Weisfeiler-Lehman (WL) kernel. We then analyze the correlation between their unsupervised WL-distance and the $L_{2}$ distance between their Gnn embeddings. The results reveal a compelling trend: as we decrease the distance threshold, we observe a strengthening correlation between WL-distance and Gnn embedding distance. This pattern supports our hypothesis, indicating that computation trees that are proximate in the WL-space also exhibit proximity in the Gnn embedding space.
@@ 3 Bonsai: Proposed Methodology
Table 2: Correlation between WL-embedding similarities and training gradients.
Dataset Correlation $p$-value
Cora 0.74 $\approx 0$
Citeseer 0.83 $\approx 0$
Pubmed 0.38 0.02
Reddit 0.42 $\approx 0$
@@
    Similar Gnn embeddings generate comparable outputs, resulting in similar impacts on the loss function and, consequently, on the gradients Table 2 presents empirical data supporting this relationship, showing statistically significant correlations ($p<0.05$) between WL-embedding similarities and training gradients.
Figure 3: Pipeline of the proposed algorithm for graph condensation using a toy example is displayed. We use $L=1$ and $k=2$ for this example.
@@
  • Representativeness: Each exemplar should be similar to a large number of other computation trees from the input training set, ensuring it captures common structural patterns. We quantify the representative power of a computation tree using the idea of reverse $k$ nearest neighbors (§ 3.1). Specifically, if tree $T_{1}$ is among the $k$-NN of tree $T_{2}$ for a small $k$, this indicates these two trees are similar in WL-embedding. Hence, we seek to include those trees in the condensed set that reside in the $k$-NN of lots of other trees. Consequently, if these trees are selected, their Gnn embeddings are also likely similar to their $k$-NN neighbors from the WL space. As a result, they can effectively approximate the Gnn embeddings of the filtered-out nodes. Since similar Gnn embeddings lead to similar gradients (Table 2), we minimize the information lost from nodes that are filtered out.
  • Diversity: The set of exemplars should be diverse, maximizing the coverage of different structural patterns present in the original graph (§ 3.2). To achieve this objective, we develop a greedy tree selection algorithm (Alg. 1), which begins with the reverse $k$-NN set and iteratively selects the computation tree that appears in the $k$-NN sets of the maximum number of currently uncovered trees. This approach ensures two key properties:
    1. 1.
      It systematically captures trees that are centrally located within the graph’s embedding space.
    2. 2.
      It progressively selects trees that provide maximum coverage of the remaining, yet-unselected computation trees.
By prioritizing trees with the highest marginal impact on uncovered trees, our method naturally creates a diverse subset that comprehensively represents the original graph’s structural and computational characteristics. The induced subgraph spanned by these exemplar trees forms the initial condensed set. This initial version undergoes further refinement through an iterative process of edge sparsification and enrichment to reduce $O(|\mathcal{E}|)$ complexity of Gnns.
@@ 3.1.1 Sampling for Scalable Computation of Reverse $k$-NN
Proof. For the formal proof, see App. A.1. Lemma 1 induces the following positive implications.
@@ 3.2 Coverage Maximization
The exemplars, $\mathbb{A}_{greedy}$, selected by Alg. 1 provides an $1-1/e$ approximation, i.e., $\Pi(\mathbb{A}_{greedy})\geq\left(1-\frac{1}{e}\right)\Pi(\mathbb{A}^{*})$.
@@
Alg. 1 begins with the reverse $k$-NN set of each computation tree as input. It then iteratively selects trees based on their marginal cardinality - specifically, choosing the tree that appears in the $k$-NN sets of the largest number of yet-uncovered trees (lines 3-4). A tree is considered uncovered if none of its $k$-nearest neighbors have been selected for the condensed set. This focus on marginal contribution naturally promotes diversity. Consider two similar trees, $T_{1}$ and $T_{2}$, both with high reverse $k$-NN cardinality. Due to the transitivity of distance functions, these trees likely share many of the same neighbors in their reverse $k$-NN sets. Consequently, if $T_{1}$ is selected, $T_{2}$’s marginal cardinality significantly decreases despite its high initial reverse $k$-NN cardinality, preventing redundant selection of similar trees.
@@
Initial condensed graph: The initial condensed graph $\mathcal{G}_{s}=(\mathcal{V}_{s},\mathcal{E}_{s},\bm{X}_{s})$ is formed by extracting the induced subgraph spanned by the exemplar computation trees, i.e., $\mathcal{V}_{s}=\left\{u\mid\exists\mathcal{T}_{v}^{L}\in\mathbb{A}_{greedy},\>u\in\mathcal{N}_{L}(v)\right\}$, $\mathcal{E}_{s}=\{(u,v)\mid u\in\mathcal{V}_{s},v\in\mathcal{V}_{s}\}$ and $\bm{X}_{s}=\{\mathbf{x}_{v})\mid v\in\mathcal{V}_{s}\}$. Set $\mathcal{N}_{L}(v)$ contains nodes within $L$ hops from $v$.
Sparsification and Enrichment: In the final step, we seek to sparsify the graph induced by the exemplar trees. Furthermore, in the additional space created due to sparsfication, we include more exemplar trees, resulting in further magnification of the representative power. Further details of our implementation is provided in App. A.4.